the area of a rectangle is represented by x^2-4. If the perimeter is 40, find x.

Answer by ewatrrr(24785)

 
Hi,
.
A = x^2-4
A = Lw = (x-2)(x-2)
P/2 = L + w = 20
(x-2) + (x+2) = 20
2x = 20
x = 10